Find the velocity of an electron with radius 10mm and magnetic field strength 0.5T


1
Expert's answer
2021-03-02T04:55:44-0500

Radius=10/1000=0.01m

V=qBr/m

=(1.6*10-19*0.01*0.5)/6.64*10-27

V=1.205*105m/s